OpenStudy (ray_masters):

Will Fan and give Medal! Which legislation justified Britain's taxing American colonists to pay for its war debt? Proclamation of 1763 Stamp Act of 1765 Declaratory Acts of 1766 Townshend Acts of 1767

OpenStudy (anonymous):

What do you think?

OpenStudy (ray_masters):

I think that the answer would be A but i'm not really sure

OpenStudy (anonymous):

No, the proclamation of 1763 was in regards to land.its a group of acts. Try again

OpenStudy (anonymous):

it was after 1765

OpenStudy (ray_masters):

I'm thinking that it's either D or B, but I'm not that good in history

OpenStudy (anonymous):

Its D.

OpenStudy (ray_masters):

Thnxs for your help

OpenStudy (anonymous):

Anytime!

OpenStudy (ray_masters):

Let me know if ya need help in any other subjects beside History,

OpenStudy (anonymous):

Are you good at Algebra?

OpenStudy (ray_masters):

A bit, do you have a problem?

OpenStudy (anonymous):

It takes 7 hours for Isabella to rake leaves by herself, but her brother Matthew can work three times as fast. If they work together how long will it take them to rake leaves? I'm given a chart with boxes for Isabella and Matthew. They are labeled: Rate Time Work Done

OpenStudy (ray_masters):

can you post a snapshot of the chart?

OpenStudy (anonymous):

Townshed acts

OpenStudy (ray_masters):

i'm sorry i don't yet understand that type of math yet @ZekeMueller and thanks @kam_Kam_356 but we already got the answer but i will fan ya anyway :)

OpenStudy (anonymous):

Can someone help me in history? I need help and thank you/ your welcome

OpenStudy (anonymous):

Im in algebra, can i give it a try

OpenStudy (anonymous):

@kam_Kam_356 What do ya need help with?

OpenStudy (anonymous):

create a scenario in which the Electoral College vote results in a tie. How would the U.S. select a president in this scenario?

OpenStudy (ray_masters):

i'm going to leave this post open so ya can help each other but i g2g

OpenStudy (anonymous):

Let's imagine a plot of 21 blocks Isabella can clear 21 blocks in seven hours, or 3 per hour. (21 / 7 = 3) Matt can clear 21 blocks in 7/3 hours, or 9 blocks per hour. (21 / 7/3 = 9) Every hour together, they finish 12 blocks (9 + 3). 21 ÷ 12 = 1 with 9 remainder, (which is 9/12 --> 3/4) The raking should take 1 hour 45 minutes.

OpenStudy (anonymous):

ok bye

OpenStudy (ray_masters):

C y'all l8ter :)

OpenStudy (anonymous):

Let's imagine a plot of 21 blocks Isabella can clear 21 blocks in seven hours, or 3 per hour. (21 / 7 = 3) Matt can clear 21 blocks in 7/3 hours, or 9 blocks per hour. (21 / 7/3 = 9) Every hour together, they finish 12 blocks (9 + 3). 21 ÷ 12 = 1 with 9 remainder, (which is 9/12 --> 3/4) The raking should take 1 hour 45 minutes. @ZekeMueller

OpenStudy (anonymous):

@kam_Kam_356 Although it was first addressed by Article II, if there were a tie in the Electoral College we would follow the process outlined in the 12th Amendment (ratified in 1804): “the House of Representatives shall choose immediately, by ballot, the President” and “the Senate shall choose the Vice-President.” The catch is that the votes in the House aren’t tallied by each representative: “the votes shall be taken by states, the representation from each state having one vote.” That means that all the representatives within a given state vote as a bloc, and each state has one vote. The majority of votes within that state bloc determine its vote. The votes in the Senate are cast by the individual senators. That also means that with an even 50 votes in the House and 50 in the Senate (D.C. doesn’t get a vote), there is potential for yet another tie vote. But the 12th Amendment covers that too: “if the House of Representatives shall not choose a President whenever the right of choice shall devolve upon them, before the fourth day of March next following, then the Vice-President shall act as President, as in case of the death or other constitutional disability of the President.” This section was later superseded by the 20th Amendment, which moved the start of the new session of Congress from March to early January. If the Senate is unable to break a tie for the vice president, according to the Presidential Succession Act of 1947, the speaker of the House serves as acting president.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Thanks @kam_Kam_356